1. You buy a stock for $50 per share and simultaneously buy a put with a $50 strike price and a $1.00 premium. If the stock price falls to $45 per share, what is your per-share net profit from this investment position? If you think net profit is negative, enter your answer using a number preceded by a minus sign. If you think net profit is positive, enter your answer as a number.

2. You buy a stock for $50 and simultaneously write a call with a strike price of $51 and a premium of $0.50 per share. If the stock price falls to $48 per share, what is your per-share net profit? If you think net profit is negative, enter your answer using a number preceded by a minus sign. If you think net profit is positive, enter your answer as a number.
Answer #1

ANSWER :


1.


Profit in dollar per share 

= Gain on bought share + Gain on put bought


= (45 - 50) + ((50 - 45) - 1)


= - 5 + 4

= - 1  (ANSWER).  (negative value indicates that there is ;loss)


2.


Profit  in dollar per share 

= Gain on bought share + Gain on call sold


= (48 - 50) + ((51 - 48) + 0.50)


= - 2 + 3.5

= 1.5  (ANSWER).

answered by: Tulsiram Garg